OpenStudy (anonymous):

I need some help with Induction. I was able to do the first part of proving them equal but after that...

OpenStudy (anonymous):

\[\sum_{K=1}^{n}=1^3+2^3+3^3...+n^3=(\frac{ n(n+1) }{ 2 })^2\]

OpenStudy (anonymous):

For Step 1 I got: Left side= \[1^3=1\] Right side= \[(\frac{ 1(1+1) }{ 2 })^2=1\]

OpenStudy (anonymous):

And every time I tried to work out the rest of the problem I didn't get anything like I was suppose too...

OpenStudy (phi):

you assume true for n, and show it is true for (n+1) the sum up to n^3 is ( n (n+1)/2)^2 now add (n+1)^3

OpenStudy (phi):

you have to show that this simplifies to ( (n+1)(n+2)/2 )^2

OpenStudy (phi):

Try putting (n+1)^3 over the common denominator of 4: \[ \frac{ (n (n+1))^2}{4}+\frac{4(n+1)^3}{4} \] or \[ \frac{ n^2 (n+1)^2+4(n+1)^3}{4} \]

OpenStudy (phi):

factor out (n+1)^2 and see what you get